A bag contains exactly three red marbles. five yellow marbles and two blue marbles.if three marbles are to be drawn from the bag at the same time,what is the probability that all three will be the same color? express your answer as a common fraction
Answer by stanbon(75887)

A bag contains exactly three red marbles. five yellow marbles and two blue marbles.if three marbles are to be drawn from the bag at the same time,what is the probability that all three will be the same color? express your answer as a common fraction
------------------
P(3red or 3yellow or 3blue) = [3C3 + 5C3 + 0]/10C3 = (1+10+0)/120 = 11/120
==========================
Cheers,
Stan H.